Lim x→1 13 √x - 7 √ x / 5√x - 3√x is :

Limx→1  13√x - 7x / 5√x - 3√x   is :

Grade:Upto college level

2 Answers

Palaash Gaunekar
13 years ago

I suppose it has to be 3  because root of 1 will not trouble you and 13-7/5-3 will be 6/2 that is 3

AskiitiansExpert Abhinav Batra
13 years ago

Dear Mainak

Limx→1  13√x - 7√x / 5√x - 3√x=13√1 - 7√1 / 5√x - 3√1

                                                         =13-7/5-3=3

ALL THE best

AskIITIANsExpert

ABhinav Batra